OpenStudy (anonymous):

If sin θ = 3/5, find cos θ?

OpenStudy (jango_in_dtown):

\[\cos \theta=\pm \sqrt{1-\sin ^{2}\theta} \]

OpenStudy (anonymous):

 You can use the Pythagorean Theorem to find the other side which will give you part of the answer to cos(theta)